# Hyperparameter Optimization for Randomized Algorithms: A Case Study for Random Features

### v2

@@ A.2 Common EKI Configuration
To train the hyperparameters in this experiment, the ensemble size was taken to be 42 for the 41 parameters #$\#\{\sigma,\{U_{ij}\}_{ij},\{D_{ii}\}_{i}\}=(1,36,4)$. The number of features used while optimizing the hyperparameters was taken to be 200. The convergence is plotted over 20 iterations and under 20 restarts of the algorithm with different random samplings from the prior in Figure 8. An experiment run with a larger noise level in the training data is shown in Figure 9.
@@
Figure 8: Convergence of the error $N_{\mathrm{ens}}^{-1}\sum_{n=1}^{N_{\mathrm{ens}}}\|\Gamma^{-\frac{1}{2}}(y_{n}-\mathcal{G}(u_{n}))\|^{2}$ for $20$ training experiments, plotted on a log-scale. The initial value is normalized to one. For this experiment, convergence stagnates at around $8$ ensemble iterations. Iterations for this plot contain $42$ ensemble members.
@@
Figure 9: Row one GP and row two RF experiments as in Figure 4, but with increased noise level of $\Sigma=10^{-2}I$.
A.5 Details for Section 4.3Accelerated Uncertainty Quantification for a Cloud Parametrization Scheme
@@
To train the hyperparameters in this experiment, the ensemble size was taken to be 635, for the 634 parameters #$\#\{\sigma,\{U_{ij}\}_{ij},\{D_{ii}\}_{i}\}=(1,630,3)$. The number of features used while optimizing the hyperparameters was taken to be 200. The convergence is plotted over 20 iterations, and under 20 restarts of the algorithm with different random samplings from the prior in Figure 10. The priors for the physical parameters in the calibration experiment (not to be confused with the hyperparameter priors) are given in Table 3.
For the MCMC algorithm, we use a simple random walk Metropolis method. The step-size was tuned in each experiment to achieve an acceptance rate of approximately 25%. The MCMC is initialized in a region of high posterior mass and the GP selected a smaller stepsize than RF one by a factor of 33~{}33. We ran the algorithm until posteriors appeared to converge, for a total of $2\times 10^{5}$ steps for RF and $5\times 10^{5}$ steps for GP.
@@
Figure 10: Convergence of the error $N_{\mathrm{ens}}^{-1}\sum_{n=1}^{N_{\mathrm{ens}}}\|\Gamma^{-\frac{1}{2}}(y_{n}-\mathcal{G}(u_{n}))\|^{2}$, normalized to be initially one over $20$ training experiments. For this experiment, convergence stagnates at around $12$ ensemble iterations. Iterations for this plot contain $635$ ensemble members.
@@
Table 3: Priors for the parameters for the EDMF calibration experiment. Priors are defined by mapping $\mathsf{N}(X,1)$ distributions into the desired bounds with a shifted and scaled logit function.
@@
static_stab_coeff $[0.01,1]$ $0.4$

### v4

@@ A.3.1 Performance Dependence on Number of Features
To tune the hyperparameters in the chaotic dynamics experiment, the EKI ensemble size was taken to be $42$ for the $41$ parameters #$\#\{\varsigma,\{U_{ij}\}_{ij},\{S_{ii}\}_{i}\}=(1,36,4)$ in the rank-$4$ nonseparable feature distribution. The number of features used while optimizing the hyperparameters was taken to be $M=200$. The convergence is plotted over $20$ iterations and under $20$ restarts of the algorithm with different random samplings from the prior in Figure 8. An experiment run with a larger noise level in the training data is shown in Figure 9; this figure should be compared with Figure 4.
@@
Figure 8: Convergence of the error $n_{\mathrm{iter}}^{-1}\sum_{n=1}^{n_{\mathrm{iter}}}\|\Gamma^{-1/2}(z_{n}-\mathcal{G}(u_{n}))\|^{2}$ for $20$ independent experiments on the Lorenz 63 system, plotted on a log-scale. The initial objective value is normalized to one. For this experiment, convergence stagnates at around $8$ ensemble iterations. Iterations for this plot contain $42$ ensemble members.
@@
Figure 9: Row one GP and row two RF experiments for the Lorenz 63 system as in Figure 4 (cf. rows two and three there), but with increased noise level corresponding to noise covariance matrix $\Sigma=10^{-2}I$.
A.4.1 Performance Dependence on Feature Distribution Structure
To illustrate the dependence on the rank under optimization, the Lorenz 63 experiment of Section 4.2 is repeated across all ranks $1,\ldots,9$ for the nonseparable feature distribution and all ranks $(1,1),\ldots,(3,3)$ for the separable feature distribution, taking all other factors (such as the number of features or optimizer options) fixed. The optimization took an mean time of 114 seconds largely independent of the rank or separability structure. In Figure 10, the resulting $L^{2}$-error of the predicted mean at each time step with both training and test data (and normalized by trajectory length) are plotted against the rank. The plot was obtained from averages over $10$ re-initializations of the optimizer. We include the performance of the GP for reference. For the nonseparable feature distribution structure, the increase in rank shows large benefits to performance on the training data until a plateau occurs at higher ranks. This motivates our choice of taking the rank equal to four in the experiments because increasing the rank quadratically increases the number of tunable parameters. The separable structure is able to achieve similar performance when the input is full rank and poor performance otherwise. Both are able to obtain performance similar to that of the GP reference, as observed in an application to forecast the attractor. We observe that in both experiments, qualitative behavior exhibited on the training trajectory is reflected in the test trajectory.
Figure 10: Results on changing the rank $r=1,\dots,9$ of the nonseparable feature distribution (left plot) covariance (48) and of the separable feature distribution covariance (right plot) with input and output ranks $r=(1,1),\ldots,(3,3)$ for the Lorenz 63 system. On each plot, the two curves represent the $L^{2}$ error of predicted time steps along the training (dashed) and the test (solid) trajectories. The GP results (independent of rank) are shown for comparison in orange alongside RF results in blue. The vertical axis shows the $L^{2}$ error of the tuned emulator against the rank on the horizontal axis.
A.5 Details for Section 4.3Uncertainty Quantification of a Cloud Model
@@ A.4.1 Performance Dependence on Feature Distribution Structure
To tune the hyperparameters in the Bayesian inverse problem experiment involving the cloud model, the ensemble size was taken to be $J=635$ for the 634 parameters given by #$\#\{\varsigma,\{U_{ij}\}_{ij},\{S_{ii}\}_{i}\}=(1,630,3)$. The number of features used while optimizing the hyperparameters was taken to be $M=200$. The convergence is plotted over 20 iterations of EKI and under 20 restarts of the algorithm with different random samplings from the prior in Figure 11. The priors for the physical parameters in the underlying geophysical inverse problem (not to be confused with the hyperparameter priors for the forward map surrogate) are given in Table 5.
@@
Figure 11: Convergence of the error $n_{\mathrm{iter}}^{-1}\sum_{n=1}^{n_{\mathrm{iter}}}\|\Gamma^{-\frac{1}{2}}(z_{n}-\mathcal{G}(u_{n}))\|^{2}$ for tuning the RF emulator used in uncertainty quantification of a cloud model, normalized to be initially one over $20$ experiments. For this experiment, convergence stagnates at around $12$ ensemble iterations. Iterations for this plot contain $635$ ensemble members.
@@
Table 5: Priors for the parameters for the EDMF calibration experiment. Priors are defined by mapping unit variance normal distributions into the desired bounds with a shifted and scaled $\mathrm{logit}$ function.
@@
static_stab_coeff $[0.01,1]$ $0.40$
For the MCMC algorithm, we use a simple random walk Metropolis method. The step size was tuned in each experiment to achieve an acceptance rate of approximately $25\%$. The MCMC method is initialized in a region of high posterior mass and the GP selected a smaller step size than the RF did by a factor of approximately three. We ran the algorithm until the posteriors appeared to converge, for a total of $2\times 10^{5}$ steps for RF and $5\times 10^{5}$ steps for GP. Table 6 compares the time taken for tuning and prediction and also displays training and test errors with different feature distributions or with GPs. The ranks chosen for the RF method were based on having low test error (as done in Appendix A.4.1). The test set was held out of the final iteration of data from the training point selection algorithm (Cleary et al., 2021); by construction, test accuracy is thus evaluated on a region of high posterior mass that the MCMC samples will explore. Table 6 shows more consistent performance on training and test error with RF, while the GP results show a much lower error on training data than on test data, indicating possible overfitting.
Table 6: Performance of the different emulators on the accelerated uncertainty quantification test problem, measured as average $L^{2}$-error over the training data and over a held-out test set, under 10 restarts of the algorithm. Reported timings are for the tuning procedure, and the average time to perform one MCMC likelihood calculation. The rank of the separable or nonseparable distribution are indicated in parentheses in column one.
Emulator Tuning Time Training Error Test Error MCMC Step Time
GP (L-BFGS-B) 260s 4.85 15.08 7.8ms
RF (untuned), rank $1$ 15.80 12.61 3.0ms (8 threads)
RF (EKI), rank $(5,1)$ 504s (8-threads) 16.77 14.05 3.8ms (8 threads)
RF (EKI), rank $1$ 565s (8-threads) 12.81 11.70 3.0ms (8 threads)
In Figure 12, the panels show the performance of RF-CES posteriors based on a separable feature distribution and on an untuned RF emulator, in addition to the emulator with nonseparable feature distribution seen in Subsection 4.3 and Figure 6 in the main text. As expected, the posterior corresponding to the untuned emulator performs poorly; its samples cover much of the geophysical parameter prior support; see Table 5. However, this is somewhat surprising in light of Table 6 because the untuned RF emulator obtained low test error there. This implies that low emulator test error for the forward map does not automatically translate to an accurate posterior for the inverse problem. The RF emulator with separable feature distribution also produces a posterior with too wide of a support compared to the GP-CES reference posterior. Finally, we observe that although the GP emulator obtained large forward map test error compared to that of our chosen nonseparable RF emulator, the resulting posterior distributions remained similar as seen in Figure 6.
Figure 12: Additional plots for the uncertainty quantification of a cloud model. The corner plots display marginal posterior distributions resulting from emulators using GP-CES (yellow), RF-CES (blue), and RF-CES without tuning (green scatter). Left: rank-$1$ nonseparable feature distribution in RF-CES, as shown in Figure 6. Right: separable feature distribution in RF-CES with rank (5,1).